A random sample is drawn from a population of known standard deviation 22.1. Construct a 95% confidence interval for the population mean based on the information given (not all of the information given need be used).

a. n=121, x¯=82.4, s=21.9 n=121, x¯=82.4, s=21.9

b. n=81, x¯=82.4, s=21.9 n=81, x¯=82.4, s=21.9


1
Expert's answer
2022-05-03T12:40:14-0400

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

a) The critical value for "\alpha = 0.05, df=n-1=120" degrees of freedom is "t_c=z_{1\u2212\u03b1\/2;n\u22121}=1.97993."


"CI=(82.4-1.97993\times\dfrac{21.9}{\sqrt{121}},"

"82.4+1.97993\times\dfrac{21.9}{\sqrt{121}})"

"=(78.458,86.342)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "78.458< \mu < 86.342," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(78.458,86.342)."


b) The critical value for "\alpha = 0.05, df=n-1=80" degrees of freedom is "t_c=z_{1\u2212\u03b1\/2;n\u22121}=1.990063."

"CI=(82.4-1.990063\times\dfrac{21.9}{\sqrt{81}},"

"82.4+1.990063\times\dfrac{21.9}{\sqrt{81}})"

"=(77.558,87.242)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "77.558< \mu < 87.242," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(77.558,87.242)."